A coupon bond that pays interest of $62 annually has a par value of $1,000, matures in 5 years, and is selling today at $75.75 discount from par value. The current yield on this bond is ____.

A) 6.20%

B) 6.71%

C)8.08%

D) 6.61%
Answer #1

Current value=1000-75.75=$924.25

Hence current yield=Annual coupon/Current value

=62/924.25

=6.71%(Approx).